# Algebraic geometry of bubbling Kähler metrics

### v1

@@ 2.2.1. Setup and algebraic construction
The following main theorems 2.4, 2.6, 2.8 aim at algebraic reconstruction and generalization of the bubbles via metrics in [DS17, Sun23]. The first version below (Theorem 2.4) avoids the new terminology of [Od24b] for the readers convenience but the next one (Theorem 2.6), which uses loc.cit, will be more general and stronger.
@@
for any $t\neq 0$. Here $\widehat{\operatorname{vol}}(-)$ denotes the normalized volume 444also called under different names in the literature such as volume density ([SS17, HS17]) up to a constant, or local volume in [XZ24] (although different from the local volume of [Ful11]) in the sense of [Li19] (as its developments are also surveyed at [LLX20], [Od24a, §2]).
@@
    The central fiber $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ of each $\mathcal{X}^{\prime}\ni\sigma^{\prime}_{0}(x)$, is an algebraic negative weight deformation ([Od24b, §2], after [CH22, Definition 1]) of a K-semistable $\mathbb{Q}$-Fano cone $(W,\xi)$ (in the sense of [CS18]) over $\mathbb{A}_{\mathbbm{k}}^{1}$ in a canonical manner, up to a finite base change, which we denote simply as $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$. In particular, $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ is K-semistable in the sense of [Sun23, §5]. The total space $\mathcal{X}^{\prime}_{\rm min}$ is achieved as an elementary modification of $\mathcal{X}$ in the sense of [HL14, AHLH23].
@@
    The central fiber $\mathcal{X}^{\prime}_{{\rm min},0}$ of $\mathcal{X}^{\prime}_{\rm min}$ which satisfies all the above conditions is unique (cf., also Theorem 2.8 for the comparison with [Sun23]).
@@
Below, we use the framework of [Od24b] more to strengthen above theorem 2.4, following the same circle of ideas. (For the readers who prefer the naive more classical Setup 1, one can skip or only skim through the following Definition 2.5, Theorem 2.6.) In particular, the total spaces $\mathcal{X}^{\prime}_{\rm min}$ can be also taken as a canonical object, and we can treat much more general degeneration families. We begin with the following additional natural terminology, to compare with Setup 1 and partially review [Od24b].
@@
  • (Canonicity) The first step construction of $\Gamma_{1,\geq 0}^{\prime}$, $\widetilde{\mathcal{X}}_{1}^{\prime}\to\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[\Gamma_{1,\geq 0}^{\prime}]]$ and its section $\sigma_{1}^{\prime}$ is canonical i.e., only need the data $\tilde{\pi}$.
@@
As a link to differential geometric works in [Sun23, dBS23], we first prove the following, which confirms [Sun23, Conjecture 5.4 (i)] and also makes a progress on Conjecture 5.4 (ii) of op.cit (cf., also [dBS23, §4.1]).
@@
to make [DS17, Sun23] works. Then,
@@
Question 2.9 (Yau-Tian-Donaldson type problem for bubbling).
Does our $\mathcal{X}^{\prime\prime}_{\rm min,0}$ for some $\mathcal{T}^{\prime\prime}$ in Theorem 2.4 coincide with the minimum bubble in the differential geometric sense of [Sun23, Definition 1.9] (Definition 2.1), obtained in [Sun23, Theorem 1.10]?
We confirmed the above question affirmatively in the following $1$ or $2$-dimensional classical cases considered in [dBS23]:
  • conical singular curves (“log curves”), considered in loc.cit §2.3,
  • $2$-dimensional $A_{k}$-singularities with the standard metric constructed by the Gibbons-Hawking ansatz, considered in loc.cit §3.3,
  • and higher dimensional $A_{1}$-singularity case, considered in loc.cit §4.2.
We omit the details here and write the proofs elsewhere (hopefully for more generality).
@@
First we discuss the construction of $\mathcal{X}^{\prime}_{\rm min}$. For simplicity, note that we can and do assume $\sigma(t)$ is the origin 0=($\vec{0}=(0,\cdots,0)$ for any $t$.
By the extended Rees construction (cf., e.g., [Mat89, Chapter 5, section 4], [Tei03, §2]), we can and do assume $v_{X}$ is associated to weights $w_{1},\cdots,w_{l}\in\mathbb{R}_{>0}$ on each coordinate of $\mathbb{A}_{\mathbbm{k}}^{l}$. See some complementary arguments in [Od24b, Example 2.9]. Below, we consider degree of polynomials on $\mathbb{A}_{k}^{l}$ with respect to this set of weights $\xi:=(w_{1},\cdots,w_{l})$. From the construction, we suppose the defining ideal of $W\subset\mathbb{A}_{\mathbbm{k}}^{l}$ is generated by the non-constant ($\xi$-)quasi-homogeneous elements $f_{1},\cdots f_{a}$ with the degrees (again with respect to $\xi$) bounded above by a positive constant $C>0$, and also that $\mathcal{X}_{t}$ for small enough $t$ is also defined by equations of degrees bounded above by $C>0$.
@@
Now we consider the set of all possible polynomials $g_{1},\cdots,g_{a}$ on $\mathbb{A}_{\mathbbm{k}}^{l}$ of the degrees at most by $C$, without constant terms, and consider the corresponding universal family $\mathcal{U}\twoheadrightarrow\mathbb{A}_{\mathbbm{k}}^{m}$ for $m\gg 0$. Then we again apply the flattening stratification and universal hull construction [Kol08, 1, 2]. We then take its $T$-invariant strata $M$ which contains the base disk of $\mathcal{X}\to\Delta$ and the base of the positive weight degeneration $X$ degenerating to $W$ of Example 2.9 of [Od24b] after [Li19], and an isotrivial degeneration of $W$ to the metric tangent cone $C_{0}(X)$ (cf., [DS17, LWX21]). One can also take $\mathcal{T}$ a (non-canonical) affine test configuration induced by an algebraic subtorus $T^{\prime}$ of $T$ with the rank $1$ (the symbol $\mathcal{T}^{\prime}$ is reserved for different family to appear later as it is more natural notation for it), as a subfamily of the whole family $\mathcal{X}_{\tau}\to U_{\tau}$ of [Od24b, Example 2.9]. For $r>1$ case, this $T^{\prime}$ is clearly non-unique as a rational approximation of the algebraic positive degeneration $X\rightsquigarrow W$ ([LX18, §3], [Od24b, §2]). We denote the base change of $\mathcal{U}\to\mathbb{A}_{k}^{m}$ to $M$ as $\mathcal{V}\to M$. Further, we can and do replace the chart $M$ of $\mathcal{M}:=[M/T]$ by its equivariant resolution of singularities, and birationally modify $\mathcal{M}$ accordingly, which we still denote by the same letters $M$ and $\mathcal{M}$ for the simplicity of notation.
@@
From our construction, the resulting $\varphi\colon[\mathcal{V}/T]\to\mathcal{M}$ is a $T$-equivariant $\mathbb{Q}$-Gorenstein family over an Artin stack which is smooth (as stack). By the openness of log terminality, we can further assume it is a log terminal morphism and all the fibers contain the origin 00.
Now we set $c:=\widehat{\operatorname{vol}}(X,0)$, the (minimized) normalized volume ([DS17, Li19], cf., also [Od24b, §2] for review) 101010recently it is also called local volume [XZ24] of the klt singularity $0\in X$ and take an open subscheme of $M$ whose $\mathbbm{k}$-valued points are $\{s\in M(\mathbbm{k})\mid\widehat{\operatorname{vol}}(0,\varphi^{-1}(s))\geq c\},$ which is indeed Zariski open by [BL21, Theorem 1], [Xu20, 1.3]. Then, we consider
$\{s\in M(\mathbbm{k})\mid\widehat{\operatorname{vol}}(0,\varphi^{-1}(s))=c\},$
which gives a Zariski closed subset $M_{c}$ of $M$ again by [BL21, Theorem 1]. We also consider reduced subscheme structure on $M_{c}$. We denote $\mathcal{M}_{c}:=[M_{c}/T]$.
@@
As $\mathcal{M}_{c}\neq\emptyset$, we consider $\varphi^{-1}(s)$ for $s\in\mathcal{M}_{c}(\mathbbm{k})$. The construction of $W$ for each such $\varphi^{-1}(s)$ gives a structure of canonical higher $\Theta$-strata in the sense of [Od24b] (following [HL14, AHLH23]), of the quotient stack $[M/T]$ supported on $[M_{c}/T]$. The first step of its proof is that, in general, if you have a log terminal affine morphism with base points $\tilde{\mathcal{X}}(\supset\sigma(S)=\{0\}\times S)\to S$ and the positive weight degenerations $\tilde{X}_{s}\ni\vec{0}\rightsquigarrow W_{s}\ni\vec{0}$ of the pointed klt affine variety $X_{s}\ni\vec{0}$ to its K-semistable cone $W_{s}$ can be taken uniformly, i.e., if there is a flat family over $S\times_{\mathbbm{k}}U_{\tau}$ with a connected algebraic $\mathbbm{k}$-scheme $S$ and an affine toric variety $U_{\tau}$ which corresponds to a rational polyhedral cone $\tau$, the corresponding minimal normalized volume is constant i.e., not depending on $s\in S$. This follows from the comparison with the leading coefficient of the index character (cf., [MSY08, Li19, CS18] and also [BB73, 4.1], [Od24a, §2]). Therefore, in particular, $\Delta\times U_{\tau}\dashrightarrow M$ which is induced by the positive weight deformation of $W$ can not be a morphism. Another technical result we use is a natural higher analogue of the $\Theta$-reductivity type statement as proved in [Che24, 1.1]. See the statements [Od24a, §3.3, Theorem 3.6] for a complementary discussion and the actual statement we use, from which it follows that $\{\mathcal{M}_{c}\}_{c}$ give a higher $\Theta$-stratification of $\mathcal{M}$, in the sense of [Od24b].

### v2

@@ 2.2.1. Setup and algebraic construction
In the following two subsubsections, we set up stages and list some statements in more precise form we prove in the later subsubsections.
@@
The following main theorems 2.4, 2.6, 2.11 aim at algebraic reconstruction and generalization of the bubbles via metrics in [DS17, Sun23]. The first version below (Theorem 2.4) avoids the new terminology of [Od24b] for the readers convenience but the next one (Theorem 2.6), which uses loc.cit, will be more general.
@@
for any $t\neq 0$. Here $\widehat{\operatorname{vol}}(-)$ denotes the normalized volume 444also called under different names in the literature such as volume density ([SS17, HS17]) up to a constant, or local volume in [XZ24] (although different from the local volume of [Ful11]) in the sense of [Li19] (see §2.1 or [LLX20], [Od24a, §2] in more details).
@@
    The central fiber $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ of each $\mathcal{X}^{\prime}\ni\sigma^{\prime}_{0}(x)$, is an algebraic negative weight deformation ([Od24b, §2], after [CH22, Definition 1]) of a K-semistable $\mathbb{Q}$-Fano cone $(W,\xi)$ (in the sense of [CS18]) over $\mathbb{A}_{\mathbbm{k}}^{1}$ in a canonical manner, up to a finite base change, which we denote simply as $\mathcal{W}\twoheadrightarrow\mathbb{A}_{\mathbbm{k}}^{1}$ or $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$. In particular, $\mathcal{X}^{\prime}_{0}:=\pi^{\prime-1}(0)$ is K-semistable in the sense of [Sun23, §5]. The total space $\mathcal{X}^{\prime}_{\rm min}$ is achieved as an elementary modification of $\mathcal{X}$ in the sense of [HL14, AHLH23].
@@
    Suppose $W=X$ and they have only isolated singularity with unobstructed deformations, and $\mathcal{X}$ is weakly non-degenerating (compare [BR15]) in the sense we reveal during the proof. For instance, if it holds for $n=2$ ADE singular case.
    Then, the central fiber $\mathcal{X}^{\prime}_{{\rm min},0}$ of $\mathcal{X}^{\prime}_{\rm min}$ can be constructed canonically (cf., also Theorem 2.11, Theorem 2.13 for partial comparison results with differential geometric bubbles [Sun23, dBS24]).
@@
Below, we use the framework of [Od24b] more to refine and generalize above Theorem 2.4 to Theorem 2.6, following the same circle of ideas. (For the readers who prefer the naive more classical Setup 1, one can skip or only skim through it.) In particular, the total space $\mathcal{X}^{\prime}_{\rm min}$ can be also taken “more canonically”, and we can treat much more general degeneration families over Novikov type rings. We prepare the following additional natural terminology, to compare with Setup 1 and partially review [Od24b].
@@
We prove both Theorems 2.4, 2.6 in the subsection §2.2.3.
@@
For the sake of organization of our explanations, here we focus and finish the issue of finiteness of bubbles, for fixed $\pi$, as a part of the previous theorem 2.4. We use [XZ24] but we note that a differential geometric counterpart is proven before for the case when $\mathcal{X}_{t}(t\neq 0)$ are all smooth, compactifiable to metrized Kähler manifolds, in [Sun23, Corollary 1.12].
Theorem 2.8.
We fix a positive integer $n$. In the setup 1, 2, 3, we modify $\mathcal{X}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$ as in Theorem 2.4. Then this modification process does not continue for infinitely many times for $\operatorname{dim}(X)=n$. Hence, we algebraically obtain a sequence of (weighted) algebraic bubbles $Z_{0},\cdots,Z_{2k}$ similarly to [Sun23, Cor 1.12].
(7) $=\widehat{\operatorname{vol}}_{\mathcal{X}_{t}}(\sigma(t)).$
for an arbitrary $0<|t|\ll 1$.
Proof.
The modification process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}_{\rm min}$ terminates by [XZ24]. We take $\mathcal{Y}$ as the end object of the repetition of the process $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime\prime}$, towards deeper bubbles. Then the last equality (6) holds as otherwise we can repeat the process by Theorem 2.4, and the equality (7) is obvious from identification of $\mathcal{X}$ and $\mathcal{Y}$ outside fibers over $0\in\Delta$. ∎
Note that the last statement of the above Theorem 2.8 is a semistable reduction type theorem corresponding to virtual compactness of “moduli” of singularities with fixed normalized volumes. In particular, the following corollary to above, of a weak version of the resolution of singularities type, is pointed out by M.Mauri.
Corollary 2.9 (Alteration of log terminal singularities).
Our main reason of writing this is on our different method than usual resolution of singularities, rather than the result itself.
Proof.
Theorem 2.10 (K-moduli of Fano cone [Od24a]).
Consider the set (resp., all $\mathbb{Q}$-Gorenstein families) of the $\mathbbm{k}$-isomorphic classes of K-polystable (resp., K-semistable) $\mathbb{Q}$-Fano cone of $n$-dimension, with the fixed normalized volume $V$ of the vertices. Then, it forms a good proper (and separated) moduli algebraic space over $\mathbbm{k}$ and its stacky enhancement, generalizing the K-moduli of $\mathbb{Q}$-Fano varieties.
2.2.2. Statements on the comparison with differential geometry
As a link to differential geometric bubblings i.e., rescaled up pointed Gromov-Hausdorff limits, as studide in [Sun23, dBS24], we first prove the following, which confirms [Sun23, Conjecture 5.4 (i)] and also makes a progress on Conjecture 5.4 (ii) of op.cit (cf., also [dBS24, §4.1]).
@@ 2.2.2. Statements on the comparison with differential geometry
to make [DS17, Sun23] works. Then, if the ambient local coordinates around $\sigma(\Delta)\subset\mathcal{X}$ are compatible with [Sun23, 4.5, 4.6],
@@
The above compatibility of local ambient coordinates mean the coincidence of (constants times) the restriction of global coordinates and $\tilde{f}^{\alpha}_{j,k}$ in [Sun23, §4, after Lemma 4.5]. We expect somewhat weaker asymptotic version is enough but anyhow the above discussions naturally motivate the following question.
Question 2.12 (Coincidence with differential geometric bubbling).
Does our $\mathcal{X}^{\prime\prime}_{\rm min,0}$ in Theorem 2.4, applied with some $\mathcal{T}^{\prime\prime}$ and good ambient local coordinates, coincide with the (differential geometric) minimum bubble in the sense of [Sun23, Definition 1.9] (Definition 2.1), obtained in [Sun23, Theorem 1.10]?
Note that we refer to a priori non-canonicity of the constructions in Theorem 2.4, which we expect to parallel the choice of the family of Kähler metrics $\{g_{t}\}_{t}$ on $\mathcal{X}\to\Delta$. Below, we confirm the above question affirmatively in the following $1$ or $2$-dimensional classical cases considered in [dBS24]. The statements are for deeper bubbles as well, due to the iterative nature.
Theorem 2.13 (Confirmation - coincidence with differential geometric bubbling).
The above question has affirmative answers at least in the following three cases, which form examples of Setup 1. More precise statements are itemized as follows. In all the cases, the semistable bubbles and polystable bubbles coincides i.e., $\mathcal{X}^{\prime}=\mathcal{X}^{\prime\prime}$ for each pair.
  1. ($\mathrm{i}$)
    (“Log curve” case) If we consider $\mathcal{X}:=\mathbb{C}\times\Delta_{t}\to\Delta_{t}$ with flat (singular) Kähler metrics with conical singularities $p_{i}(t)\in\mathbb{C}$ for $i=0,\cdots,k$ of angle $2\pi(1-\beta_{i})$ with $10<\beta_{i}<$ 1, the sequence of bubblings constructed algebro-geometrically by Theorem 2.4 (cf., Remark 2.7) can be taken canonically, coincides with that of [dBS24, §2.2, Lemma2] and hence coincides with differential geometric bubbles i.e., rescaled pointed Gromov-Hausdorff limits (or, polarized limit space in the sense of [DS17], to be more precise).
  2. ($\mathrm{ii}$)
    (ALE case with $n=2$) Consider general smoothing affine family of ADE singularity $\mathcal{X}(\supset\Delta_{t})\to\Delta_{t}$ where $X=\mathcal{X}_{0}$ has only one singularity at $\sigma(0)=0$ which is of $A_{k}$ type. Suppose we have a continuous family of ALE hyperKähler (orbi-)metrics $g_{t}$ on $\mathcal{X}_{t}$ whose Kähler forms $(\omega_{I,t},\omega_{J,t},\omega_{K,t})$ ($I$ is the original complex structure for $X_{t}$s in Setup1) satisfy that
    • $\omega_{I,t}$ is $\partial\overline{\partial}$-exact
    • $\omega_{J,t}+\sqrt{-1}\omega_{K,t}$ is holomorphic on $\mathcal{X}\setminus X$.
    Then, the sequence of bubblings constructed algebro-geometrically by Theorem 2.4 can be taken canonically again, coincides with the construction of [dBS24, §3.3, Theorem 3] and hence coincides with differential geometric bubbles.
  3. ($\mathrm{iii}$)
    ($A_{1}$ case with any $n$) If $\mathcal{X}\to\Delta_{t}$ is open subset of proper holomorphic family over $\Delta$ which is the smoothing of higher dimensional $A_{1}$-singularity, the minimal bubbling constructed algebro-geometrically by Theorem 2.4 is $V(z_{1}^{2}+\cdots+z_{n}^{2}-1)\subset\mathbb{A}^{n+1}$ and coincides with the differential geometric minimal bubble with the Stenzel asymptotic conical metric ([EH79, Ste93]).
We proceed to the proofs of the above Theorems 2.4, 2.6, 2.11, and 2.13.
2.2.3. Details of the algebraic algorithmic construction
In this subsubsection, we prove Theorem 2.4 and 2.6.
@@ 2.2.3. Details of the algebraic algorithmic construction
First we discuss the construction of $\mathcal{X}^{\prime}_{\rm min}$, its central fiber $\mathcal{X}^{\prime}_{{\rm min},0}$ and its degeneration to a K-semistable klt cone $W\curvearrowleft T$ (recall at §2.1). We consider the deformations of $W$ and here we a priori face the problem of infinite-dimensionality of the deformation space of affine varieties, especially those with non-isolated singularities (see [SJZ23, §6, 6.9]). Nevertheless, following the idea of negative weight deformation ([CH22], also cf., [Sun23, dBS24]), we can circumvent the problem.
First, we take a temporary embedding of $W$ to $\mathbb{A}^{l}$ via (non-constant) $T$-semiinvariants $z_{1},\cdots,z_{l}$ so that the vertex maps to the origin. Suppose its Reeb vector field $\xi$ which corresponds to the volume minimizer and has weights $w_{1},\cdots,w_{l}\in\mathbb{R}_{>0}$ on each coordinate of $\mathbb{A}_{\mathbbm{k}}^{l}$ i.e., ${\rm deg}_{\xi}(z_{i})=w_{i}$ (cf. [DS17, Od24a, Od24b]). We consider finite generators $f_{1},\cdots,f_{N}$ of the defining ideal which gives $W\subset\mathbb{A}_{\mathbbm{k}}^{l}$ and set $d_{i}:={\rm deg}_{\xi}(f_{i})$. Consider all affine deformations of $W$ of the $\xi$-negative type $V(\{f_{i}+h_{i}\}_{i=1,\cdots,N}\mid 0<{\rm deg}_{\xi}(h_{i})<{\rm deg}_{\xi}(f_{i}))$. Also compare more intrinsic but conditional related construction: Corollary 2.17. Then, by applying the flattening stratification and taking the universal hull [Kol08, 1, 2], we obtain a $T$-equivariant affine flat $\mathbb{Q}$-Gorenstein (locally stable) family $\pi_{\mathcal{Y}}\colon\mathcal{Y}\to{\rm Def}^{-}(W)\subset\mathbb{A}^{m}$ for $m\gg 0$. We consider its $T$-invariant locally closed subset
$Z^{+}:=\{b\in{\rm Def}^{-}(W)\mid\widehat{\rm vol}(0\in\pi_{\mathcal{Y}}^{-1}(b))=\widehat{\rm vol}(0\in W))\},$
by [BL21]. By replacing ${\rm Def}^{-}(W)$ by a $T$-invariant open affine neighborhood by Sumihiro’s theorem ([Sum74]), we can suppose $Z^{+}$ is its closed subset. Further, by [Che24], $\mathcal{O}_{\pi_{\mathcal{Y}}^{-1}(Z^{+})}$ has ideal sequences $\{I_{\lambda_{i}}\}_{\lambda_{i}\in S(W)\subset\mathbb{R}_{>0}}$ such that for any $b\in Z^{+}$$\{I_{\lambda_{i}}|_{\pi_{\mathcal{Y}}^{-1}(b)}\}_{\lambda_{i}\in S(W)\subset\mathbb{R}_{>0}}$ gives the degeneration to the K-semistable Fano cone à la [DS17]. Here, $S(W)$ is the holomorphic spectrum of $W$ also in the sense of [DS17]. Recall these at §2.1.
Now, we take a positive integer $m$ and shrink ${\rm Def}^{-}(W)(\ni[W])$ enough again to a $T$-invariant open affine neighborhood such that all $i=1,\cdots,m$, $I_{\lambda_{i}}/I_{\lambda_{i+1}}$ are all free $\mathcal{O}_{Z^{+}}$-modules of rank $r_{i}$, and have $T$-equivariant splittings of $I_{\lambda_{i}}\twoheadrightarrow I_{\lambda_{i}}/I_{\lambda_{i+1}}\simeq\mathcal{O}_{Z^{+}}^{r_{i}}$ which corresponds to $T$-semiinvariant $\{z_{i,j}\in I_{\lambda_{i}}\mid 1\leq i\leq m,1\leq j\leq r_{i}\}$ ([IS15, 2.5]). By the complete reductivity of $T$, one can also extend $z_{i,j}$ to $T$-semiinvariant regular functions on whole $\mathcal{Y}$, which vanishes at the 00-section. With large enough $m$ and shrunk enough ($T$-invariant) $[W]\in{\rm Def}^{-}(W)$, we can and do assume the above $z_{i,j}$ generates $\Gamma(\mathcal{O}_{\pi_{\mathcal{Y}}^{-1}(b)})$ as $\mathbbm{k}$-algebras for any $b\in{\rm Def}^{-}(W)$. From the construction, it follows that for any $b\in Z^{+}$, the natural $T$-action with the weights $w_{i}$ give the degeneration to K-semistable Fano cone. We here reset ł$\l:=\sum_{i=1}^{m}r_{i}$ and $\xi=(w_{1},\cdots,w_{1},\cdots,w_{m},\cdots,w_{m})$. Then, the above $z_{i,j}$ gives a large reembedding of $\mathcal{Y}\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}\times{\rm Def}^{-}(W)$.
Next, we enlarge ${\rm Def}^{-}(W)$ a little just to include $\mathcal{X}_{t}$s. Suppose the reembedding $W\subset\mathbb{A}_{\mathbbm{k}}^{l}$ constructed above is defined by $c$ polynomial equations $g_{1},\cdots,g_{c}$. Then, take an arbitrary extended embedding of $\mathcal{X}$ to $\mathbb{A}_{\mathbbm{k}}^{l}\times_{\mathbbm{k}}\Delta$ and suppose that $\mathcal{X}_{t}$ for generic $t$ is defined by equations of $\xi$-degrees bounded above by some $C>0$. Now we consider the set of all possible polynomials $g_{1},\cdots,g_{c}$ on $\mathbb{A}_{\mathbbm{k}}^{l}$ of the degrees at most by $C$, without constant terms, and consider the corresponding universal family $\mathcal{U}\twoheadrightarrow\mathbb{A}_{\mathbbm{k}}^{p}$ for $p\gg 0$. There is a corresponding morphism $\iota\colon\Delta\to\mathbb{A}_{\mathbbm{k}}^{p}$ to this $\mathcal{X}\to\Delta$.
Then we set $M:={\rm Def}^{-}(W)\cup T\cdot\iota(\Delta)\subset\mathbb{A}_{\mathbbm{k}}^{p}$, possibly after shriking $\Delta$ if necessary. Recall that $\Delta$ in Setup 1 (b) is a pointed smooth affine $\mathbbm{k}$-curve so that $M$ underlies a structure of algebraic $\mathbbm{k}$-scheme. Here, this $T$-action on $\iota(\Delta)\subset\mathbb{A}_{\mathbbm{k}}^{p}$ is given via the re-embedding coordinates and is the same for the whole $T$-action on $M$.
@@
We denote the base change of $\mathcal{U}\to\mathbb{A}_{k}^{p}$ to $M$ as $\mathcal{V}\to M$ and $\mathcal{M}:=[M/T]$. (For those who prefer rank $1$ setup, one could alternatively take $\mathcal{T}$ a (non-canonical) affine test configurations induced by an algebraic subtorus $T^{\prime}$ of $T$ with the rank $1$ (the symbol $\mathcal{T}^{\prime}$ is reserved for different family to appear later as it is more natural notation for it), as a subfamily of the whole family $\mathcal{X}_{\tau}\to U_{\tau}$ of [Od24b, Example 2.9]. For $r>1$ case, this $T^{\prime}$ is clearly non-unique as a rational approximation of the algebraic positive degeneration $X\rightsquigarrow W$ ([LX18, §3], [Od24b, §2]).)
@@
From our construction, the resulting $\varphi\colon[\mathcal{V}/T]\to\mathcal{M}$ is a $\mathbb{Q}$-Gorenstein log terminal family over an Artin stack and all the fibers contain the origin 00.
@@
Now it follows that $[Z^{+}/T]$ has a structure of canonical higher $\Theta$-strata ([Od24b]) on $\mathcal{M}$ near $[W]$ by the above construction, which is also locally liftable. This follows from the above construction of the $T$-acted base, using [Che24], together with the following fact: if you have a log terminal affine morphism with base points $f\colon\tilde{\mathcal{X}}(\supset\sigma(S)=\{0\}\times S)\to S$ and a uniform positive weight degeneration $\tilde{X}_{s}\ni\vec{0}\rightsquigarrow W_{s}\ni\vec{0}$ of the pointed klt affine variety $X_{s}\ni\vec{0}$ to its K-semistable cone $W_{s}$, i.e., if there is a flat family over $S\times_{\mathbbm{k}}U_{\tau}$ with a connected algebraic $\mathbbm{k}$-scheme $S$ and an affine toric variety $U_{\tau}$ which corresponds to a rational polyhedral cone $\tau$, the corresponding minimal normalized volume $\widehat{\rm vol}(f^{-1}(s)\ni 0)$ is constant i.e., not depending on $s\in S$. This follows from the comparison with the leading coefficient of the index character (cf., [MSY08, Li19, CS18] and also [BB73, 4.1], [Od24a, §2]). On the other hand, over a neighborhood of $[W]\in Z^{+}$, our new $T$-action gives a higher $\Theta$-strata by [Che24].
Hence, we can apply semistable reduction type theorem [Od24b, Theorem 3.8 (also cf., 1.1)] (which generalizes [Lan75, AHLH23]) and we obtain a Langton type modification of the original morphism $f\colon\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}\to[M/T]$ (resp., $f\colon\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to[M/T]$) which corresponds to $\mathcal{X}$. For Setup 1 (b) case, recalling the arguments in §3 of [Od24b] and its notations, we consider $\overline{f}$ and its restriction $f_{\rm new,\xi}\colon{\rm Spec}(R_{\rm new}(\xi^{\prime}))\to\mathcal{M}$ as in its Corollary 3.10. It corresponds to an affine faithfully flat family over $\mathbbm{k}[[M^{\prime}_{\xi^{\prime}_{\geq 0}}]]$. Since $R_{\rm new}(\xi^{\prime})=\mathbbm{k}[[M^{\prime}_{\xi^{\prime}_{\geq 0}}]]$, we can consider its arbitrary restriction in the sense of Lemma 3.5 of §3.1 of [Od24b] and denote it as $\mathcal{X}^{\prime}_{\rm min}$ (see the statements of Theorem 2.4). For the analytic variant Setup 1 (a) case, it is easy to see that by the above procedure and a restriction in an analogous sense to Lemma 3.5 of op.cit, one obtains $\mathcal{X}^{\prime}_{\rm min}\to\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$  [^10] Here, $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ denotes the discrete valuation ring which consists of the convergent power series of $t$. possibly after replacement of $t$ by its $d$-th root for some $d\in\mathbb{Z}_{>0}$. We are now going to confirm the required conditions of $\mathcal{X}^{\prime}_{\rm min}$ in our Theorem 2.4 and that $\mathcal{W}$ of ($\mathrm{iv}$) also naturally show up during the construction.
As noted in [Od24b], note that the family $\mathcal{X}^{\prime}_{\rm min}$ (as the total space) is non-canonical which corresponds to a choice of $T^{\prime}$ while the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ does not depend on it. Nevertheless, under certain conditions, they can be constructed canonically. For instance, suppose $W$ has only isolated singularity and have unobstructed deformation i.e., with smooth $T$-equivariant deformation space $[W]\in{\rm Def}(W)$ which is algebraized by [Art69, Pin74]. Then, by [KR82, BH85, LP17], $[W]\in{\rm Def}(W)$ is $T$-equivariantly isomorphic to $0\in\mathbb{A}_{\mathbbm{k}}^{d}$ with linear $T$-action. This gives a $T$-equivariant decomposition ${\rm Def}(W)=\oplus_{i\in I}V_{i}$ to $T$-eigenspaces $V_{i}$. We denote the $\xi$-weights on $V_{i}$ as $W_{i}$. Now we add an assumption that $\iota\colon\Delta\to{\rm Def}(W)$ which corresponds to $\mathcal{X}$ has non-vanishing differential at each each $V_{i}$: we call this condition weakly non-degenerating, comparing with stronger conditions in [BR15]. A subtle point is that because of a priori non-universality of ${\rm Def}(W)$, $\iota$ is not canonical and only its differential $\partial\iota_{t=0}$ is canonical by the versality. On the other hand, the construction of $\mathcal{X}^{\prime}_{\rm min}$ shows that they are determined by $\partial_{t}\iota|_{V_{i}}$ for those $V_{i}$ with $W_{i}=1$. Thus, the weak non-degeneracy assumption implies that the above construction of $\mathcal{X}^{\prime}_{\rm min,0}$ is canonical. This confirms ($\mathrm{vi}$).
The obtained morphism from $\mathop{\mathrm{Spec}}\nolimits\mathbb{C}[[t]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$), or its finite extension $\mathbb{C}[[t^{\frac{1}{d}}]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t^{\frac{1}{d}}]]$) for some $d\in\mathbb{Z}_{>0}$, to $\mathcal{M}$ gives the desired family $\mathcal{X}^{\prime}_{\rm min}\to\Delta$ from the construction.